# Guiding Large Language Models to Post-Edit Machine Translation with Error Annotations

### v1

@@ 4.1 Datasets
[…]
@@ 4.2 Metrics
[…]
@@ 5 Prompting Results
[…]
@@
[…]
@@
[…]
@@
[…]
@@
FT (Bi) ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
FT (Multi) ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
@@
FT (Bi) ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
FT (Multi) ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
@@
FT (Bi) ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
FT (Multi) ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
Table 4: Fine-tuning (FT) results for LLaMA-2 7B. prompt ($k$=0) and prompt ($k$=10) indicate the zero- and 10-shot prompting results of LLaMA-2 7B respectively. FT (Generic): Fine-tuning with generic feedback; FT (Bi): Fine-tuning with fine-grained feedback in bilingual setting, where models are individually fine-tuned for each language pair; FT (Multi): Fine-tuning with fine-grained feedback in multilingual setting, combine 3 language pairs to fine-tune a single model. We test the statistically significance of improvements over the best prompting baseline and $\dagger$ marks results with $p$-value $\leq 0.05$.
@@ 6.1 Automatic Evaluation
[…]
@@ D.2 LLaMA-2 13B Results
prompt ($k$=0) 0.50 0.74 0.71
prompt ($k$=10) 0.50 0.61 0.73
FT (Bi) ${}^{\dagger}$ 0.61 ${}^{\dagger}$
FT (Multi) ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
@@
prompt ($k$=0) 0.51 0.68 0.75
prompt ($k$=10) 0.58 0.56 0.79
FT (Bi) 0.57 ${}^{\dagger}$ ${}^{\dagger}$
FT (Multi) ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
@@
FT (Multi) ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
Table 7: Fine-tuning (FT) results for LLaMA-2 13B. prompt ($k$=0) and prompt ($k$=10) indicate the zero- and 10-shot prompting results of LLaMA-2 13B respectively. FT (Bi): Fine-tuning in bilingual setting, where models are individually fine-tuned for each language pair; FT (Multi): Fine-tuning in multilingual setting, combine 3 language pairs to fine-tune a single model. We test the statistically significance of improvements over the best prompting baseline and $\dagger$ marks results with $p$-value $\leq 0.05$.
@@ F.3 Feedback from Annotators
Zh-En 0 0.47 / 0.75 / 0.70 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
10 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
En-De 0 0.45 / 0.86 / 0.69 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
10 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
En-Ru 0 0.43 / 0.82 / 0.73 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
10 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / 0.70$\ddagger$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
[…]
Original Generic Score MQM InstructScore xCOMET
Zh-En 0 0.47 / 0.75 / 0.70 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / 0.72 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
10 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
En-De 0 0.45 / 0.86 / 0.69 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / 0.72 - ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
10 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
En-Ru 0 0.43 / 0.82 / 0.73 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / 0.80 / 0.73 - ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
10 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
@@
Zh-En 0 0.47 / 0.75 / 0.70 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / 0.70 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
10 0.48 / ${}^{\dagger}$ / ${}^{\dagger}$ 0.48 / ${}^{\dagger}$ / ${}^{\dagger}$ 0.48 / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / 0.63 / ${}^{\dagger}$ 0.50 / ${}^{\dagger}$ / ${}^{\dagger}$
En-De 0 0.45 / 0.86 / 0.69 0.42 / ${}^{\dagger}$ / ${}^{\ddagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
10 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
En-Ru 0 0.43 / 0.82 / 0.73 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\ddagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / 0.73 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\ddagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
10 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
@@
Zh-En 0 0.47 / 0.75 / 0.70 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / 0.70 0.44 / ${}^{\dagger}$ / ${}^{\dagger}$ 0.47 / ${}^{\dagger}$ / ${}^{\dagger}$
10 0.48 / 0.65 / ${}^{\dagger}$ ${}^{\ddagger}$ / 0.64 / ${}^{\dagger}$ ${}^{\dagger}$ / 0.62 / ${}^{\dagger}$ ${}^{\dagger}$ / 0.64 / ${}^{\dagger}$ 0.51 / 0.62 / ${}^{\dagger}$
En-De 0 0.45 / 0.86 / 0.69 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / 0.72 - ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
10 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
En-Ru 0 0.43 / 0.82 / 0.73 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$
10 ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$ ${}^{\dagger}$ / ${}^{\dagger}$ / ${}^{\dagger}$

### v2

@@ 4.1 Datasets
Figure 2: Zero- and 10-shot prompting results for LLaMA-2 7B. Top: BLEU scores for Chinese-English (Zh-En), English-German (En-De), English-Russian (En-Ru) pairs; Middle: Translation Edit Rate (TER) where zero-shot results show the amount increased compared to that of 10-shot; Bottom: COMETDA scores. Note that we only report the supporting language pair (zh-en) results for InstructScore. Numerical results for both 7B and 13B are in Appendix E.1.
@@ 4.2 Metrics
We report the traditional BLEU metric Papineni et al. (2002) with exponential smoothing as implemented in the sacrebleu toolkit Post (2018), the Translation Edit Rate (TER) Snover et al. (2006) which is the minimum number of edits needed to change a hypothesis so that it exactly matches one of the references, normalized by the average length of the references, and a modern neural metric, the reference-based COMETDA score Rei et al. (2020). Scores for all these metrics are reported in the 0-1 range.
@@ 5 Prompting Results
For all language pairs, we observe a marginal improvement when post-editing with any form of feedback in zero-shot settings, with small increases in BLEU COMETDA scores, along with reduced TER. Although the score differences between the original and post-edited MT can be small, they are statistically significant ($p\leq 0.05$) for all cases. One exception is Zh-En pair, for which BLEU drops by 0.01 to 0.02 points after integrating feedback other than MQM.
@@
The improvements from zero to 10-shot prompting are shown by hashed lines in Figure 2. The performance gap between the original and post-edited MT widens with few-shot learning. We examine a consistent gain in both BLEU and COMETDA scores, which represent the overall MT quality. The average gain across language pairs is +0.04 BLEU (on a 0-1 scale) and +0.03 for COMETDA. TER, which measures the remaining amount of edits to be made also shows -0.03 point improvement for Zh-En, -0.06 point for En-De, and -0.04 point for En-Ru.
@@
The 13B model unsurprisingly achieve higher BLEU and COMETDA and lower TER compared to the 7B model in zero-shot settings. However, this performance gap narrows down with the increase in number of few-shot examples. This trend suggests that few-shot learning helps bridge the performance gap between model sizes for MT post-editing. We report comprehensive results on LLaMA-2 13B in Appendix E.
@@
For 10-shot prompting, most forms of our tested feedback, regardless of granularity, converge to a similar performance. However, while the two MT quality metrics, BLEU and COMETDA remains similar for different forms of feedback, there is a clear difference for TER. When providing generic feedback or automatic annotations from xCOMET, TER decreases by approximately 0.15 points for Zh-En and 0.3 points for En-De and En-Ru compared to the original baseline. Score-based feedback remains to show the least increase in performance, but they also decrease 0.1 points for Zh-En and 0.2 points for En-De and En-Ru, which are statistically significant. Nevertheless, prompting does not reveal a marked advantage for using certain type of feedback for post-editing.
@@
FT (Bi) 0.53 0.63 0.76
FT (Multi) 0.53 0.61 0.76
@@
FT (Bi) 0.56 0.58 0.79
FT (Multi) 0.59 0.55 0.79
@@
FT (Bi) 0.51 0.65 0.80
FT (Multi) 0.52 0.63 0.80
Table 4: Fine-tuning (FT) results for LLaMA-2 7B. prompt (kk=0) and prompt (kk=10) indicate the zero- and 10-shot prompting results of LLaMA-2 7B respectively. FT (Generic): Fine-tuning with generic feedback; FT (Bi): Fine-tuning with fine-grained feedback in bilingual setting, where models are individually fine-tuned for each language pair; FT (Multi): Fine-tuning with fine-grained feedback in multilingual setting, combine 3 language pairs to fine-tune a single model. We test the statistically significance of improvements over the best prompting baseline and $\dagger$ marks results with $p$-value $\leq 0.05$.
@@ 6.1 Automatic Evaluation
We compare the fine-tuning results of each language pair against the original translation quality (indicated as ‘Original’ in Table 4). Across language pairs, metrics of MT quality all increase for fine-tuning. Translation quality increases steeply by approximately +0.07 BLEU, +0.08 COMETDA and -0.21 TER on average for all language pairs. The multilingual approach mostly outperforms the bilingual one, suggesting that the advantages gleaned from fine-tuning with diverse language pairs outweigh the benefits of matching the fine-tuning data language consistent to the test language pair. We observe the same trend with LLaMA-2 13B in Appendix Table 7: fine-tuning results improve upon the original baseline results by +0.1 BLEU, +0.08 COMETDA and -0.25 TER points on average.
@@ D.2 LLaMA-2 13B Results
prompt (kk=0) 0.50 0.74 0.71
prompt (kk=10) 0.50 0.61 0.73
FT (Bi) 0.51 0.61 0.77
FT (Multi) 0.54 0.58 0.76
@@
prompt (kk=0) 0.51 0.68 0.75
prompt (kk=10) 0.58 0.56 0.79
FT (Bi) 0.57 0.55 0.80
FT (Multi) 0.60 0.53 0.80
@@
FT (Multi) 0.54 0.56 0.81
Table 7: Fine-tuning (FT) results for LLaMA-2 13B. prompt (kk=0) and prompt (kk=10) indicate the zero- and 10-shot prompting results of LLaMA-2 13B respectively. FT (Bi): Fine-tuning in bilingual setting, where models are individually fine-tuned for each language pair; FT (Multi): Fine-tuning in multilingual setting, combine 3 language pairs to fine-tune a single model. We test the statistically significance of improvements over the best prompting baseline and $\dagger$ marks results with $p$-value $\leq 0.05$.
@@ F.3 Feedback from Annotators
Language Shots BLEU (↑) / TER (↓) / COMETDA (↑)
Original Generic Score MQM InstructScore xCOMET
Zh-En 0 0.47 / 0.75 / 0.70 0.45 / 0.63 / 0.71 0.46 / 0.69 / 0.7 0.48 / 0.72 / 0.70 0.45 / 0.63 / 0.72 0.45 / 0.62 / 0.72
10 0.51 / 0.63 / 0.73 0.51 / 0.64 / 0.73 0.51 / 0.65 / 0.72 0.50 / 0.61 / 0.74 0.51 / 0.60 / 0.75
En-De 0 0.45 / 0.86 / 0.69 0.55 / 0.57 / 0.78 0.51 / 0.68 / 0.75 0.51 / 0.68 / 0.75 - 0.57 / 0.54 / 0.80
10 0.58 / 0.55 / 0.79 0.55 / 0.59 / 0.80 0.56 / 0.58 / 0.79 - 0.53 / 0.57 / 0.81
En-Ru 0 0.43 / 0.82 / 0.73 0.51 / 0.60 / 0.77 0.48 / 0.70 / 0.74 0.48 / 0.69 / 0.74 - 0.51 / 0.58 / 0.80
@@
Zh-En 0 0.47 / 0.75 / 0.70 0.50 / 0.66 / 0.73 0.50 / 0.72 / 0.72 0.50 / 0.74 / 0.71 0.50 / 0.61 / 0.75 0.53 / 0.59 / 0.76
10 0.51 / 0.61 / 0.74 0.51 / 0.61 / 0.74 0.50 / 0.61 / 0.73 0.50 / 0.61 / 0.75 0.54 / 0.58 / 0.76
En-De 0 0.45 / 0.86 / 0.69 0.58 / 0.55 / 0.80 0.49 / 0.73 / 0.73 0.48 / 0.76 / 0.72 - 0.60 / 0.52 / 0.81
10 0.58 / 0.54 / 0.80 0.58 / 0.55 / 0.80 0.58 / 0.54 / 0.80 - 0.62 / 0.51 / 0.82
En-Ru 0 0.43 / 0.82 / 0.73 0.53 / 0.56 / 0.79 0.46 / 0.74 / 0.74 0.44 / 0.80 / 0.73 - 0.55 / 0.54 / 0.83
10 0.54 / 0.55 / 0.80 0.54 / 0.56 / 0.80 0.53 / 0.56 / 0.80 - 0.57 / 0.52 / 0.85
@@
Zh-En 0 0.47 / 0.75 / 0.70 0.43 / 0.69 / 0.73 0.45 / 0.66 / 0.74 0.40 / 0.70 / 0.70 0.42 / 0.68 / 0.73 0.41 / 0.67 / 0.73
10 0.48 / 0.65 / 0.74 0.48 / 0.64 / 0.74 0.48 / 0.64 / 0.73 0.48 / 0.63 / 0.75 0.50 / 0.63 / 0.76
En-De 0 0.45 / 0.86 / 0.69 0.42 / 0.68 / 0.73 0.51 / 0.63 / 0.76 0.54 / 0.62 / 0.76 - 0.50 / 0.6 / 0.77
10 0.54 / 0.60 / 0.78 0.54 / 0.60 / 0.78 0.54 / 0.59 / 0.78 - 0.56 / 0.57 / 0.79
En-Ru 0 0.43 / 0.82 / 0.73 0.40 / 0.72 / 0.72 0.45 / 0.66 / 0.73 0.48 / 0.64 / 0.74 - 0.45 / 0.63 / 0.76
10 0.50 / 0.61 / 0.78 0.49 / 0.62 / 0.77 0.50 / 0.61 / 0.77 - 0.52 / 0.59 / 0.81
@@
Zh-En 0 0.47 / 0.75 / 0.70 0.40 / 0.71 / 0.73 0.45 / 0.69 / 0.72 0.47 / 0.72 / 0.70 0.44 / 0.65 / 0.75 0.47 / 0.64 / 0.76
10 0.48 / 0.65 / 0.74 0.48 / 0.64 / 0.74 0.49 / 0.62 / 0.74 0.48 / 0.64 / 0.75 0.51 / 0.62 / 0.76
En-De 0 0.45 / 0.86 / 0.69 0.53 / 0.59 / 0.78 0.50 / 0.71 / 0.74 0.48 / 0.74 / 0.72 - 0.57 / 0.54 / 0.80
10 0.53 / 0.60 / 0.78 0.55 / 0.58 / 0.70 0.55 / 0.58 / 0.79 - 0.55 / 0.58 / 0.80
En-Ru 0 0.43 / 0.82 / 0.73 0.47 / 0.61 / 0.77 0.47 / 0.72 / 0.74 0.47 / 0.72 / 0.74 - 0.51 / 0.58 / 0.81
10 0.47 / 0.62 / 0.77 0.49 / 0.60 / 0.79 0.49 / 0.60 / 0.78 - 0.49 / 0.62 / 0.81